Operations Report – November 13th

Crew 198 Operations Report 13-Nov-2018

SOL: 3

Name of person filing report: Dana Levin

Non-nominal systems: Report filing, Rover Demos, engineering airlock electrical sockets, engineering airlock light, lower deck hab wall heater,

Notes on non-nominal systems:

– Generator update; after resetting the inverters to enable it to communicate load requirements to the generator generator seems to be responding to load needs and functioned without difficulty last night. It was started at 83% today and appears to be functioning well as of this report

– Report filing repeating from last night for broadcasting: we have had continued trouble with non updated report templates requiring a back and forth between MDRS and mission support demanding reports filed in different formats. This yielded significant crew frustration

o Proposed solution: reports could be turned into google forms. The links to these forms sent to the station crew to standardize filing format. This would also allow centralized data collection and presentation including some values that could be graphed over time and organized for presentation/analysis easily.

o We attempted to formulate such forms but our internet connection did not allow connection to the google servers

– Sojourner: Shannon had used Sojourner two days ago but the battery drained to the point where she was unable to safely return to the habitat. Rover was left about 4 miles north of the Hab beside brae highway.

o Our schedule was rearranged to enable 2 of our instructors to leave the hab to recover the rover in the Hab car while our students were undergoing lectures by a third instructor. This was completed without incident and luckily our pre-planned schedule was sufficient to absorb this unexpected non crew/non sim related contingency.

– Engineering air lock light and power sockets are still inoperable.

– Demos Rover remains with battery drained, water has been replaced and rover is hooked up to charge cable, will evaluate in morning

– 350 ATV is inoperable, likely with a dead battery that has yet to be replaced

Generator (hours run): 12

Solar— SOC % (Before generator is run at night): 83% (see above)

Diesel Reading – 60%

Propane Reading – not taken

Ethanol Free Gasoline – 13 gallons.

Water (auxillary tank) – 64 gallons

Water (static tank) – 350 gallons

Auxillary to Static tank transfer– yes

Gallons transferred: 30

Water in GreenHab – tank at 90%, unsure how many gallons this is

Water (loft) – Static to Loft Pump used – yes

Water Meter: unknown where this is located

Toilet tank emptied: yes

Deimos rover used: no


Beginning charge: 0%

Ending charge:

Currently charging: yes

Sojourner rover used: no


Beginning charge: 100%

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: 1

Beginning charge: 100

Ending charge: 85

Currently charging: yes

Opportunity rover used: yes

Hours: 1

Beginning charge: 100

Ending charge: 87

Currently charging: yes

Curiosity rover used:

Hours: 1

Beginning charge: 100

Ending charge: 81

Currently charging: yes

Notes on rovers: (See above)

Demos: batteries refilled, attempting recharging overnight

Sojourner recovered and appears functional

ATV’s Used: (Honda, , 350.1, 350.2)

Reason for use: EVA

Oil Added? no

ATV Fuel Used: 1 Gals

# Hours the ATVs were Used today: 1

Notes on ATVs: one of the 350s is still non functional. The battery appears dead

HabCar used and why, where? No

CrewCar used and why, where? no

General notes and comments:

Please standardize reports and consider loading them into google forms or some kind of email template with a hab specific email. This saves data by preventing redownloading of mail to different accounts, avoids incorrect forms transmitted to crew or filed by crew, and makes data transfer easier

Summary of internet: functional
Summary of suits and radios: using our own
Summary of Hab operations: nominal despite facility difficulties Summary of GreenHab operations: none
Summary of ScienceDome operations: none
Summary of RAMM operations: none
Summary of any observatory issues: none identified as we are not using observatory
Summary of health and safety issues: none

Questions, concerns and requests to Mission Support: see above

Copyright © The Mars Society. All rights reserved. | Main Site